The measure of each interior each interior angle of a regular polygon is 36 more than its adjacent exterior angle. How many sides has the polygon? Thanks!
Answer by stanbon(75887)

The measure of each interior angle of a regular polygon is 36 more than its adjacent exterior angle. How many sides has the polygon? Thanks!
---------------------------
The two angles are supplementary.
So interior = exterior + 36
x = (180-x)+36
x = 90+18 = 108
Therefore the exterior angle is 180-108 = 72
---------------------------
The sum of the exterior angles is 360
# of exterior angles is 360/72 = 5
---
So the # of sides is 5.
==================================
Cheers,
Stan H.